how many grams O2 to burn 15.0 gal C8H18

2C8H18+25O2=16CO2+18H2O
Answer #1

Before starting the calculation we need to know the density of Octane (C8H18) which is 0.703 g/mL.

Now start by converting the volume of octane from gallons to ml.

(15.0 gallons) x (100ml/0.0264gallons) = 56818.18182 ml of octane

Now find the grams of octane by dividing the volume of octane by the density.

(56818.18182 ml) x (0.703g/1ml) = 39943.18182 grams of octane

Now convert grams of octane to moles by divding by the gram molar mass, 114.23 g/mole.

(39943.18182 grams) / (114.23 g /mol) = 349.6733066 moles of octane

Now use the reaction ration (25 moles O2/2 mole of C8H18) from the balanced equaiton to convert to moles of O2.

(349.6733066 mol C8H18) x (25 moles O2/2 mole C8H18) = 4370.916 moles of O2 are needed.

Last step convert the moles of O2 needed to complete the reation to grams by multiplying by the gram molar mass of O2, 32 g/mole.

(4370.916 mole O2) x (32 g/mole) = 139869.32 grams of O2 are needed to complete the reaction



answered by: raleigh